OpenStudy (anonymous):

A cylinder has dimensions r=5cm, h=10cm. Find the approximate increase in volume wehen r increases by 0.2 cm and h decreases by 0.1 cm.

OpenStudy (imstuck):

The volume formula for a cylinder is this:\[V=\pi r ^{2}h\]So if your original values are r = 5 and h = 10, then the volume for that particular cylinder is this:\[V=\pi (5)^{2}(10)\]which is equal to \[V=250\pi \]

OpenStudy (imstuck):

When the values for r and h change to 5.2 and 9.9 respectively, the new volume is this:\[V=\pi (5.2)^{2}(9.9)\]which gives you \[V=267.7\pi \]

OpenStudy (imstuck):

So the volume increases by\[267.7\pi -250\pi =17.7\pi \]

OpenStudy (anonymous):

thank you very much sir but my options are 53.98cm3 50.98cm3 54.18cm3 54.98cm3